OpenStudy (anonymous):

a man travels a distance of 1.5 meter towards the east, then 2.0 meter towards the south and finally 4.5m towards the east in 4 s. Calculate (2)distance travelled and (b) displacement of the body

OpenStudy (anonymous):

displacement = 6.3245 m

OpenStudy (anonymous):

distance travelled = 1.5 m + 2.0 m + 4.5 m = 8.0 m

OpenStudy (anonymous):

so wats d use of time ? is that useless?

OpenStudy (anonymous):

I think it is just given to confuse... It plays no part in distance calculation nor displacement calculation...

OpenStudy (anonymous):

ya may be ...

OpenStudy (anonymous):

displacement will be equal to the hypotenuse of the Right triangle having base 6m (1.5 + 4.5) altitude 2m Using Pythagoras theorem Hypotnuse^2 = (6m)^2 + (2m)^2 = 36m^2 + 4m^2 = 40m^2 so Hypotenuse = root(40m^2) = 2root(10) m or 6.325 m

OpenStudy (anonymous):

we can also use the graph method starting form origin (0,0) and the final position of the man ,i.e, (6,2) From Distance formula D=sqroot* ( (6-0)^2 +(2-0)^2 ) D= 6.3245 m

OpenStudy (anonymous):

hey guys if u want to make a ²(square sign) then press alt+253

OpenStudy (anonymous):

thanks sukarang !!!